### v1

@@ Abstract
While large language models (LLMs) excel at handling long-context sequences, they require substantial key-value (KV) caches to store contextual information, which can heavily burden computational efficiency and memory usage. Previous efforts to compress these KV caches primarily focused on reducing memory demands but were limited in enhancing latency. To address this issue, we introduce FastKV, a KV cache compression method designed to enhance latency for long-context sequences. To enhance processing speeds while maintaining accuracy, FastKV adopts a novel Token-Selective Propagation (TSP) approach that retains the full context information in the initial layers of LLMs and selectively propagates only a portion of this information in deeper layers even in the prefill stage. Additionally, FastKV incorporates grouped-query attention (GQA)-aware KV cache compression to exploit the advantages of GQA in both memory and computational efficiency. Our experimental results show that FastKV achieves 2.00$\times$ and 1.40$\times$ improvements in time-to-first-token (TTFT) and throughput, respectively, compared to HeadKV, the state-of-the-art KV cache compression method. Moreover, FastKV successfully maintains accuracy on long-context benchmarks at levels comparable to the baselines. Our code is available at https://github.com/dongwonjo/FastKV.
@@ 1 Introduction
Figure 1: Comparison of accuracy, TTFT, throughput across different KV cache compression methods on LLaMA-3.1-8B-Instruct
Recent advancements in large language models (LLMs) have enabled the processing of long-context sequences, such as those comprising 128k tokens Achiam et al. (2023); Gemini Team et al. (2023); Anthropic . This capability significantly broadens the range of applications for LLMs Yi et al. (2024); Laban et al. (2023); Gu (2023). However, as the length of the input sequence increases, the size of these key-value (KV) caches also increases, making them a significant bottleneck in the serving of long-context LLMs. Therefore, compression of KV caches is essential for optimizing the operation of LLMs.
There has been active research on compressing KV cache to alleviate its burden in long-context handling Zhang et al. (2023); Xiao et al. (2023; 2024); Oren et al. (2024); Chen et al. (2024). Some techniques have demonstrated the ability to preserve accuracy comparable to full-context processing after the compression Li et al. (2024); Feng et al. (2024); Fu et al. (2024), while there exists a compression technique that has achieved inference speedup in both the prefill and generation stages Shi et al. (2024). However, no KV cache compression techniques currently exist that can simultaneously preserve accuracy and achieve inference speedup, particularly during the prefill stage.
In this paper, we propose FastKV, an innovative approach designed to expedite long-context handling with LLMs while maintaining their accuracy. FastKV stems from our finding that the properties of attention maps differ between the early and later layers of LLMs. Through detailed analysis of LLMs, we discover that in the later layers, attention maps concentrate on a limited set of significant tokens, which remain consistent across layers. In contrast, the early layers engage with a broader array of tokens and exhibit diverse attention patterns as they process the context of the input prmopt. Based on these findings, FastKV adopts a novel Token-Selective Propagation (TSP) method that applies different strategies of KV cache compression at the early and later layers respectively As shown in Figure 1, this dual-strategy approach enables FastKV to improve both time-to-first-token (TTFT) and throughput of long-context processing while effectively preserves the accuracy of LLMs. Our experimental results show that FastKV can achieve 2.00$\times$ speedup for TTFT and 1.40$\times$ improvement for throughput compared to HeadKV, while maintaining a similar level of accuracy with less than a 1% accuracy gap. These results demonstrate that FastKV promises a practical solution for long context scenarios, particularly in real-time applications that require efficient KV cache management and low latency for the prefill and generation stages.
@@ 2.1 Long-Context Processing with LLMs
LLMs have significantly enhanced the natural language processing (NLP) abilities of AI systems Brown et al. (2020); Radford (2018); Kamalloo et al. (2023); Jiang et al. (2021) through the use of attention mechanisms Vaswani (2017). These attention mechanisms carefully evaluate the relationships between tokens within sentences to construct context vectors by aggregating token data according to the attention scores to extract context-specific information Brauwers & Frasincar (2021). Distinct from earlier NLP models such as RNNs Sherstinsky (2020) and LSTMs Hochreiter (1997), attention mechanisms consider the entire token data from input sequences, thereby avoiding issues of prompt forgetting. Benefiting from these advancements, recent researches indicates that LLMs are capable of processing long-context sequences.
@@
However, the very nature of LLMs, which retain all token data to facilitate attention mechanisms, leads to significant memory overhead. Typically, token data for attention mechanisms are stored in the form of KV cache, and the size of KV cache escalates with sequence length, becoming a primary source of inefficiency in LLMs when processing long-context scenarios. For instance, in the LLaMA-3.1-8B model Dubey et al. (2024), managing a sequence of 128k tokens requires 17.12GB of memory for KV cache, whereas the model itself only requires 16.06GB. Therefore, the excessive KV cache demands form a critical bottleneck in LLM inference during long-context processing, and efficient compression of KV cache is crucial for streamlined LLM operation.

### v2

@@ Abstract
While large language models (LLMs) excel at handling long-context sequences, they require substantial key-value (KV) caches to store contextual information, which can heavily burden computational efficiency and memory usage. Previous efforts to compress these KV caches primarily focused on reducing memory demands but were limited in enhancing latency. To address this issue, we introduce FastKV, a KV cache compression method designed to reduce latency for long-context inference. FastKV improves processing speed while preserving accuracy by adopting Token-Selective Propagation (TSP). This approach preserves full-context information in early layers of LLMs and selectively propagates only a portion of this information in later layers. This design enables FastKV to minimize redundant computation without sacrificing contextual fidelity. Our experimental results show that FastKV achieves up to 1.97$\times$ and 4.82$\times$ improvements in time-to-first-token (TTFT) and throughput, respectively, compared to baseline without KV cache compression. Moreover, FastKV successfully maintains accuracy within 1% of the baseline on long-context benchmarks. Our code is available at https://github.com/dongwonjo/FastKV.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have enabled the processing of long-context sequences, such as those comprising 128K tokens [4, 5, 6, 9, 23]. This capability significantly broadens the range of applications for LLMs [31, 32, 33]. However, as the length of the input sequence increases, the size of these key-value (KV) caches also increases, making them a significant bottleneck in the serving of long-context LLMs. Therefore, compression of KV caches is essential for optimizing the operation of LLMs.
There has been active research on compressing KV cache to alleviate the burden in long-context handling [10, 11, 24, 25, 26, 12]. Some techniques have demonstrated the ability to preserve accuracy comparable to full-context processing after the compression [13, 14, 15], while there exists a compression technique that has achieved inference speedup in both the prefill and generation stages [16]. However, no existing KV cache compression techniques can simultaneously preserve accuracy and achieve inference speedup, particularly during the prefill stage.
To bridge this gap, we propose FastKV, a KV cache compression framework that accelerates long-context processing without compromising accuracy. FastKV adopts a novel Token-Selective Propagation (TSP) method that applies different strategies of KV cache compression at the early and later layers respectively. TSP enables early layers to attend to the full input context while propagating only a small set of critical tokens to later layers. As illustrated in Figure 1(a), this dual-strategy enables FastKV to improve both time-to-first-token (TTFT) and throughput of long-context processing while effectively preserving the accuracy of LLMs. Our experiments show that FastKV delivers improvements of up to 1.97$\times$ in TTFT and 4.82$\times$ in throughput with less than a 1% accuracy drop compared to baselines. These results demonstrate that FastKV promises a practical solution for long-context scenarios, particularly in real-time applications that require efficient KV cache management and low latency for the prefill and generation stages.
Figure 1: (a) Comparison of LongBench accuracy, TTFT, throughput across different KV cache compression methods on LLaMA-3.1-8B-UltraLong-1M-Instruct. (b) End-to-end latency breakdown measured on a 128K input tokens with 512-token generation. SnapKV [13] represents typical previous works that compress KV cache for the generation stage only.
@@ 2.1 Long-Context Processing with LLMs
LLMs have significantly enhanced the natural language processing (NLP) abilities of AI systems [27, 28, 29, 30] through the use of attention mechanisms [21]. However, the very nature of LLMs, which retain all token data to facilitate attention mechanisms, leads to significant memory overhead. Typically, token data for attention mechanisms are stored in the form of KV cache, and the size of KV cache escalates with sequence length, becoming a primary source of inefficiency in LLMs when processing long-context scenarios. For instance, in the LLaMA-3.1-8B model [7], managing a sequence of 128K tokens requires 17.18GB of memory for KV cache, whereas the model itself only requires 16.06GB. These excessive KV cache demands form a critical bottleneck in LLM inference during long-context processing, and efficient compression of KV cache is crucial for streamlined LLM operation.
Moreover, to fully realize acceleration in long-context processing, KV cache compression algorithms must be carefully designed. Previous works [13, 14, 15] have mainly focused on compressing the KV cache for the generation stage only, and this approach fails to reduce prefill latency (Figure 1(b)). Therefore, to enable efficient long-context processing, it is crucial to develop methods that not only compress the KV cache but also accelerate the computation in the prefill stage.
2.2 KV Cache Compression
KV cache compression techniques for long-context processing are primarily designed to prune tokens, thereby reducing the associated KV caches and effectively managing memory usage. A common approach leverages the attention scores generated by each attention head to selectively prune tokens in an attention-head-wise manner. This method operates on the premise that token importance varies across attention heads. A notable example is SnapKV in Figure 2(a). It discovers that each attention head in the model consistently focuses on specific prompt token during generation, and this robust pattern can be obtained from an observation window located at the end of the input prompt. Thus, SnapKV evaluates the token importance score for each attention head by summing the attention scores assigned to each token as follows:
Here, $S_{i}^{l,h}$ is the importance score of $i$-th token in $h$-th attention head of the $l$-th layer. $Att_{l}$ denotes the attention score matrix of $l$-th layer, while $N_{I}$ and $N_{obs}$ indicate the number of tokens in the input prompt and the observation window, respectively. SnapKV applies average pooling on the scores to prevent sparse token selection. Here, $12w_{p}+$ 1 indicates the size of pooling window. SnapKV selects tokens from the observation window first and then selects tokens with the highest importance scores for KV caching. Since SnapKV relies on the attention scores produced by each attention head, it requires propagating the full input prompt throughout the entire model, during the prefill stage. As a result, SnapKV cannot improve computational efficiency or reduce the latency of the prefill stage, which directly impacts TTFT, a key factor in optimizing user experience for LLM serving.
AdaKV and HeadKV in Figure 2(b)(c) extend SnapKV by applying dynamic KV budget allocation to head groups [22] and individual heads, respectively, which improves accuracy under extreme compression such as when the average KV budget is 128. However, when applied to more realistic KV budget targets, such as 512, 1024 and 2048, dynamic KV budget allocation exhibits marginal accuracy improvements or even a slight accuracy degradation compared to SnapKV, while they introduce notable latency and throughput overhead. The primary reason of the overhead is that the processing speed of the attention mechanism is limited by the attention head with the longest KV cache. As the size of KV budget increases, the imbalance in KV cache sizes across attention heads becomes more pronounced and it amplifies latency overhead. This overhead makes these methods less practical for real-world applications, where achieving low latency is critical.
In contrast, GemFilter in Figure 2(d) adopts a fundamentally different approach that enhances the computational efficiency of the prefill stage. This method identifies relevant tokens at the layer level in intermediate layers, called the filter layer, and uses these indices to compress the input prompt. Then, the prefill stage is recomputed with compressed input prompt. By processing full-context with part of the LLM, GemFilter effectively reduces TTFT. However, as the removal of tokens from the input prompt results in the complete loss of any information embedded in the discarded tokens, GemFilter struggles to maintain general performance on complex tasks or tasks requiring comprehensive contextual understanding, even when given moderate KV budgets such as 2K. Thus, there are currently no KV cache compression techniques that can adequately preserve accuracy while simultaneously improving TTFT and throughput. This underscores the need for new methods that effectively balance these objectives.
3 Proposed FastKV
3.1 Overview of FastKV
Figure 3: Illustration of the proposed FastKV scheme. The proposed FastKV introduce Token-Selective Propagation approach to selectively propagate only a limited set of tokens while effectively compressing KV data.
GemFilter accelerates the prefill stage while performing effective KV cache compression. However, by removing a portion of the tokens from the input prompt, it inevitably causes complete loss of the discarded information. Critically, LLMs rely on attention mechanisms to enable rich interactions across tokens, and GemFilter fails to exploit this characteristic. For instance, GemFilter performs well on multi-document QA tasks in the LongBench evaluation, which primarily require information retrieval, but exhibits substantially lower performance on tasks requiring comprehensive context understanding, such as code completion. To overcome this limitation, we propose FastKV, a KV cache compression framework that accelerates the prefill stage while preserving global context.
The overview of FastKV is illustrated in Figure 3. To speed up the prefill stage, FastKV employs the Token-Selective Propagation (TSP) designed to maintain full-context processing in early layers and selectively propagates only a small set of semantically important tokens to later layers. The TSP method identifies a subset of crucial tokens (e.g., 2K) at a mid-layer (TSP layer) and propagates only these tokens through subsequent layers. These selected tokens, having already aggregated contextual information through the attention mechanisms in the early layers, encapsulate the essential semantics of the input sequence and thus preserve global context throughout the later layers. Furthermore, the TSP approach reduces the number of tokens processed in the later layers, thereby accelerating the processing speed in these layers.
To perform KV cache compression, FastKV evaluates the attention scores generated by each attention head group to determine the importance of KV data. Based on the head-wise scores $S_{i}^{j,h}$ (Equation 1), the group-level importance is computed as follows:
$S_{i}^{l,g}=\frac{1}{H_{G}}\sum_{h=h_{g}}^{h_{g}+H_{G}}S_{i}^{j,h}$ (2)
Here, $H_{G}$ denotes the number of attention heads in each group. $g$ and $h_{g}$ denotes the group index and the index of the first head in the group. While FastKV applies the same KV cache compression technique to both early and later layers, in the later layers, compression is performed over a reduced set of KV data corresponding to the tokens selected by the TSP approach, allowing for more precise identification of important information. It is important to note that KV data compression occurs solely during the storing of KV data to establish KV caches for input prompt. Consequently, FastKV ensures that the entire set of KV data, including those of the removed tokens, is propagated to the subsequent layer until the TSP layer. This approach guarantees full-context propagation to preserve essential contextual information from the input prompt. This dual strategy enables FastKV to achieve significant speed improvements without compromising accuracy, as it properly preserves the contextual information of input sequences.
3.2 Token-Selective Propagation
TSP improves prefill efficiency by propagating only a subset of important tokens while retaining critical contextual information. It is defined by three factors: the layer for token selection (TSP layer), identification of important tokens and the number of tokens propagated (TSP length).
At the selected TSP layer, token importance is assessed by averaging attention scores across all heads as follows:
$S_{i}^{TSPlayer}=\frac{1}{H}\sum_{h=0}^{H-1}S_{i}^{TSPlayer,h}$ (3)
Here, $H$ denotes the number of attention heads in each layer. Since compressing the layer output requires a single index set, we average the attention weights to calculate a score $S_{i}^{TSPlayer}$ that represents the importance of tokens at the layer level, as outlined in Equation 3. The top-k tokens based on $S_{i}^{TSPlayer}$ are selected, and their indices are combined with window tokens to form the final set of TSP tokens indices propagated to the next layer. These tokens are used to index hidden states after the MLP block, effectively reducing the layer input size for subsequent computation.
Unlike GemFilter, which restarts computation from the first layer using only selected tokens, TSP continues inference from the TSP layer without recomputation. Moreover, whereas prior methods retain hidden states for all tokens through the prefill stage, TSP discards low-importance tokens after the TSP layer, reducing computation while preserving contextual information. To validate this design, we further analyze the effect of decoupling TSP from the KV budget in Section 3.3 and how TSP maintains global semantics despite this token reduction in Section 3.4.
3.3 TSP Length and KV Budget
Figure 4: (a) Minimum match rate between the 2K important tokens selected at each layer from 128K tokens and those selected by its subsequent layers. (b) Comparison of normalized L2 distances between hidden states generated by the full-context baseline, FastKV, and GemFilter-like methods. The GemFilter-like method represents the approach where the input prompt is pruned by retaining only selected tokens and inference is recomputed, following the strategy of GemFilter.
As shown in Figure 4(a), we measure the minimum match rate as the lowest overlap ratio between the top-2K important tokens selected at a given layer from 128K input tokens and those selected by all subsequent layers. The match rate of the important tokens is around 25%. This rate is sufficiently high to justify the use of the TSP approach with an adequate TSP length (e.g. 2K). However, a short TSP length (e.g. 512) may lead to significant information loss. Therefore, TSP length, which defines the number of tokens selected for propagation, is strategically designed to operate independently of the KV cache budget. For instance, if the KV cache budget is set to 512 while the TSP length is 2048, TSP propagates 2048 tokens during its operation, but only the final 512 tokens are stored in the KV cache. This separation allows TSP to process a significantly larger set of tokens during propagation, enabling it to capture more contextual information while maintaining the KV cache within its predefined budget. As shown in Section 4.4, a TSP length of 2048 has been found sufficient to preserve accuracy of the model under specific KV budget while maintaining efficiency in terms of TTFT.
3.4 Impact of TSP on LLM Output
While FastKV is designed to accelerate the prefill stage while preserving the model’s ability to process inputs in full-context, it is critical to empirically verify whether TSP effectively retains global contextual information. In this section, we quantitatively evaluate the impact of TSP on the distribution of LLM hidden states compared to a GemFilter-like approach.
As shown in Figure 4(b), we use a sampled needle-in-a-haystack data for 128K input prompt and identify the top-2K important tokens at a target layer based on attention scores computed via Equation 3. We perform inference over the 128K input prompt for the full-context baseline, indexing the output hidden states corresponding to the selected tokens. For the GemFilter-like variant, we prune the input prompt by keeping only the selected tokens and recompute inference. For the FastKV, we apply TSP at the target layer, propagate only the selected tokens through subsequent layers. We then measure the L2 distance between the hidden states obtained from the full-context inference and those obtained from FastKV and GemFilter, respectively, using the Mean Squared Error (MSE) metric. Figure 4(b) shows the normalized L2 distances across layers. For 2 different models, FastKV consistently achieves substantially lower L2 distances compared to the GemFilter-like variant, particularly from the middle layers onward. This indicates that TSP effectively preserves the global structure of hidden states, maintaining a close approximation to the full-context model despite aggressive token reduction. In contrast, the GemFilter-like approach exhibits consistently large L2 distances across layers, suggesting that prompt-level token pruning causes an irreversible loss of contextual information. These findings highlight the advantage of applying token selection after early-layer attention mixing, as in TSP, which enables the model to retain essential semantic information across the sequence. Thus, FastKV successfully balances prefill acceleration and contextual fidelity, achieving compression without sacrificing the model’s ability to capture global context.
3.5 TSP Layer Selection
Building on the analysis in Section 3.4, we describe the method for selecting an appropriate TSP layer to balance the trade-off between acceleration and performance in LLM inference. Unlike GemFilter, which selects a layer based on its ability to attend to specific tokens (e.g., a "needle" token), FastKV selects the TSP layer based on how effectively the compressed computation at that layer preserves the output distribution of the original full-context model. Specifically, we compute the MSE between the hidden states of the full-context model $\mathbf{H}$ and those of the TSP-compressed model $\mathbf{H^{\prime}_{L}}$ at various candidate layers $L$, and choose the layer that minimizes this discrepancy, as defined by the following objective:
$L_{TSP}=\operatorname*{argmin}_{L\leq L_{\max}}\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{H}_{i}-\mathbf{H}^{\prime}_{L,i}\right|_{2}^{2}$ (4)
Here, $N$ is the number of tokens used for TSP indices and $L_{max}$ denotes the maximum allowable TSP depth, set to half the total number of layers, to ensure sufficient opportunity for acceleration without applying TSP to overly later layers. This strategy ensures that TSP is applied at a point where the model has already integrated sufficient global context, enabling downstream layers to benefit from compression without degrading semantic fidelity.
4 Experiments
Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Synthetic
Method NrtvQA Qasper MF-en HotpotQA 2WikiMQA MuSiQue GovReport QMSum MultiNews TREC TriviaQA SAMSum LCC RB-P PCount PRe Avg.
LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = Full
Full KV 23.59 39.33 51.80 56.89 45.53 34.66 33.57 22.96 26.10 73.50 90.88 42.49 62.03 67.05 2.00 89.50 47.62
LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = 512
SnapKV 24.53 36.47 50.26 55.91 45.49 34.98 26.67 22.69 24.34 73.00 90.05 41.57 60.61 64.01 2.00 89.00 46.35
AdaKV 23.79 36.56 50.77 56.37 45.08 34.48 26.97 22.93 24.19 73.00 90.80 42.01 61.32 65.83 2.00 88.50 46.54
HeadKV 23.76 37.88 51.19 55.65 45.64 34.93 29.73 22.94 26.01 73.00 90.71 40.28 62.54 66.19 2.00 88.50 46.93
GemFilter 16.44 26.31 46.58 47.38 42.32 27.87 26.21 19.42 22.33 70.00 78.17 37.20 29.98 38.39 2.00 44.00 35.91
FastKV 23.34 36.35 51.84 56.29 46.44 34.11 26.13 22.83 24.10 74.00 90.50 41.24 60.98 65.22 2.00 87.50 46.43
LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = 2048
SnapKV 22.86 38.14 52.41 56.49 45.98 34.82 31.46 22.89 25.97 73.50 91.00 42.18 62.52 67.12 2.00 89.50 47.43
AdaKV 22.97 38.67 52.36 56.46 44.98 34.63 31.04 23.22 25.80 73.50 90.99 42.45 62.12 67.49 2.00 89.50 47.39
HeadKV 22.47 38.84 52.18 55.62 45.14 34.84 32.20 23.10 26.08 73.50 90.99 41.80 62.65 67.02 2.00 89.00 47.34
GemFilter 21.22 35.02 53.58 56.37 47.02 31.29 31.36 20.72 25.62 73.50 89.86 39.82 48.54 52.56 1.50 92.00 45.00
FastKV 23.65 38.49 52.98 56.80 46.83 34.23 30.83 23.05 25.76 75.50 90.99 42.72 62.00 67.14 2.00 87.50 47.53
Mistral-Nemo-12B-Instruct, KV Budget = Full
Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 1.50 98.00 48.51
Mistral-Nemo-12B-Instruct, KV Budget = 512
SnapKV 23.58 40.12 55.17 47.91 45.69 25.17 23.77 22.48 24.08 74.00 89.44 43.09 68.31 62.15 1.00 97.50 46.47
AdaKV 24.19 39.65 54.30 48.16 45.43 26.60 23.06 22.85 23.31 74.00 89.26 43.09 67.38 62.41 1.50 98.50 46.48
HeadKV 25.88 40.57 57.12 48.11 46.11 25.26 26.19 22.92 25.25 74.50 89.74 41.81 68.79 64.76 1.00 97.50 47.22
GemFilter 25.45 37.64 53.91 52.83 51.26 33.02 26.23 19.42 23.87 65.50 84.16 40.05 38.03 41.13 4.00 57.50 40.88
FastKV 25.51 41.48 56.76 48.73 45.43 26.54 23.58 22.20 24.00 74.00 89.42 43.22 70.23 65.55 1.00 98.00 47.23
Mistral-Nemo-12B-Instruct, KV Budget = 2048
SnapKV 24.35 42.95 56.92 49.08 46.43 26.26 28.27 23.68 26.07 75.00 89.82 44.20 68.89 68.00 1.00 97.50 48.03
AdaKV 24.67 42.46 56.05 48.60 45.97 26.23 26.58 23.39 25.74 75.00 89.66 44.09 68.25 67.70 1.50 97.00 47.68
HeadKV 26.05 42.15 57.46 48.49 45.85 25.86 30.27 24.22 26.17 75.00 89.66 43.74 68.64 67.29 1.50 98.50 48.18
GemFilter 26.42 42.40 56.98 57.64 53.92 34.33 30.51 21.60 25.96 72.00 89.65 44.48 48.34 48.06 3.50 97.00 47.05
FastKV 26.61 43.37 57.12 49.41 47.22 26.27 28.40 23.27 25.81 75.00 89.72 45.04 70.28 69.24 0.50 98.00 48.53
4.1 Setup
Models and Datasets. We evaluate two open-source LLMs of different sizes: LLaMA-3.1-8B-UltraLong-1M-Instruct [8] and Mistral-Nemo-12B-Instruct [20]. Built on LLaMA-3.1-8B-Instruct [7], LLaMA-3.1-8B-UltraLong-1M-Instruct is a model trained to support 1M context window size with 32 decoder layers. Mistral-Nemo-12B-Instruct has 40 decoder layers and employs a context window size of 128K tokens. To assess long-context understanding and retrieval capabilities of the LLMs, we use two benchmark datasets: LongBench [17] and Needle-in-a-Haystack [18]. See Appendix B for more details.
Implementation Details. We integrate our proposed FastKV method upon self-attention implementation of HuggingFace Transformers library, which utilizes FlashAttention-2 [19] kernel. We select layer 15 as the TSP layer for LLaMA-3.1-8B-UltraLong-1M-Instruct and layer 19 for Mistral-Nemo-12B-Instruct. We fix the observation window size ($N_{obs}$) to 8 and the kernel size of average pooling ($12w_{p}+$ 1) to 7. The TSP lengths are set to 2048 for evaluation.
Baselines. We compare our approach with four baseline methods for KV cache compression: SnapKV [13], AdaKV [14], HeadKV [15], and GemFilter [16]. To ensure a fair comparison, SnapKV, AdaKV, and HeadKV use the same local window size, pooling method, and pooling kernel size as FastKV. The indices of GemFilter filter layers for LLaMA-3.1-8B-UltraLong-1M-Instruct and Mistral-Nemo-12B-Instruct are 13 and 19, respectively [16].
4.2 Accuracy Evaluation
LongBench. The accuracy evaluation results on LongBench are summarized in Table 1. On LLaMA-3.1-8B-UltraLong-1M-Instruct, previous works such as SnapKV, AdaKV, and HeadKV, which employ token removal after prefilling all the tokens at each layer, successfully maintain accuracy after KV cache compression, with an average accuracy drop of less than 0.69% and 0.19% compared to Full KV for 512 and 2048 KV budgets, respectively. GemFilter, which recomputes the prefill stage with the fragmented input prompt which consists only of selected tokens, completely loses the information of the discarded tokens. This leads to a significantly larger accuracy drop, with reductions of up to 11.71% and 2.62% compared to Full KV for 512 and 2048 KV budgets, respectively. On the other hand, the proposed FastKV preserves accuracy on par with Full KV and other methods such as AdaKV and HeadKV. Moreover, on Mistral-Nemo-12B-Instruct, FastKV outperforms other methods and consistently achieves strong accuracy across LongBench, while achieving memory and computational efficiency comparable to GemFilter. The advantage of FastKV is especially pronounced in code completion tasks, which demand a comprehensive understanding of global context.
Figure 5: Needle-in-a-Haystack results of LLaMA-3.1-8B-UltraLong-1M-Instruct with 2048 KV budget.
Figure 6: Latency and throughput evaluation results of LLaMA-3.1-8B-UltraLong-1M-Instruct on a single NVIDIA A100 GPU. KV budgets and TSP length are both set to 2048 throughout the experiments. (a) TTFT with context length of 128K and 256K tokens. (b) Token throughput with context length 128K.
Needle-in-a-Haystack. We present the Needle-in-a-Haystack (NIAH) evaluation results for LLaMA-3.1-8B-UltraLong-1M-Instruct with a KV budget of 2048 in Figure 5. The evaluation is conducted with input context lengths up to 256K tokens. Additional results for other LLMs and KV budgets are provided in Appendix C.2. Notably, FastKV achieves performance comparable to other methods even under 256K long-context inputs. Taken together with the LongBench evaluation, this results demonstrate that FastKV effectively maintains general-purpose LLM performance across diverse long-context tasks while achieving efficient KV cache compression and inference acceleration.
4.3 Efficiency Evaluation
We evaluate the time-to-first-token (TTFT) of the prefill stage, throughput of token generation and memory usage of KV cache with LLaMA-3.1-8B-UltraLong-1M-Instruct on a single NVIDIA A100 GPU. Figure 6 reports TTFT results with 128K/256K input, and throughput results with 128K input and 2048 KV budget. More detailed evaluation results can be found in Appendix C.3.
TTFT. As shown in Figure 6(a), Both GemFilter and FastKV achieve significant TTFT improvements over other baselines by reducing the computational complexity of the prefill stage through the concept of propagating only selected tokens during this stage. As a result, FastKV achieves TTFT that is 1.97$\times$ shorter than Full KV with 128K input and 1.99$\times$ shorter than HeadKV with 256K input. While both methods enhance efficiency, GemFilter incurs additional overhead due to the need for prefill recomputation, whereas FastKV introduces extra processes for compressing generated KV caches. These two factors offset each other. Consequently, GemFilter’s slight advantage in TTFT is primarily attributed to its earlier filter layer (13) compared to the TSP layer (15) for LLaMA-3.1-8B-UltraLong-1M-Instruct. In contrast, SnapKV, AdaKV, and HeadKV exhibit even longer TTFT than Full KV. This is because they process full-context information during the prefill stage, similar to Full KV, while also incurring additional overhead from indexing and selecting crucial tokens.
Throughput. We measure the throughput for generating a total of 128 tokens. As shown in Figure 6(b), all KV compression techniques result in throughput improvements compared to Full KV. Both FastKV and GemFilter achieve a significant level of throughput improvement. FastKV’s throughput reaches 4.82$\times$ of Full KV and 1.97$\times$ of HeadKV with 2048 KV budget. In contrast, SnapKV, which stores KV caches of different tokens set for each attention head, shows slightly lower throughput. Additionally, HeadKV, which allocates varying KV budgets across attention heads, exhibits the lowest throughput among KV cache compression techniques. Managing different sequence lengths for keys and values across attention heads increases the complexity of the attention mechanism. AdaKV also adopts fine-grained KV budget allocation similar to HeadKV except that its granularity is attention-group. AdaKV mitigates the influence of fragmented KV budget to some extent thanks to GQA-aware compression, but it does not achieve the same level of throughput with FastKV and GemFilter.
Figure 7: Comparison across different KV cache compression methods
Memory Usage. As shown in Table 2, FastKV, GemFilter, and AdaKV reduce KV cache memory usage by 31.81$\times$ over the Full KV baseline for 256K tokens by retaining only a selected subset of tokens. This is achieved by maintaining shared token indices per KV head group, unlike SnapKV and HeadKV which store separate indices for each attention head.
A comprehensive comparison of latency, throughput, and accuracy across all methods is summarized in Figure 7, highlighting FastKV as the only approach that achieves favorable performance across all metrics.
Table 2: KV cache memory usage for a 256K tokens in LLaMA-3.1-8B-UltraLong-1M-Instruct. KV budget is set to 2048 for each compression method.
Method Full KV SnapKV AdaKV HeadKV GemFilter FastKV
KV Cache Size (GB) 34.36 4.30 1.08 4.30 1.08 1.08
Figure 8: LongBench results for LLaMA-3.1-8B-UltraLong-1M-Instruct. (a) Effect of TSP layer index on accuracy and TTFT (KV budget = 2048). (b) Effect of TSP length across KV budgets.
4.4 Ablation Studies
TSP Layer Index. We evaluate the impact of the TSP layer index on model accuracy under a fixed KV budget of 2048. As shown in Figure 8(a), applying TSP at earlier layers degrades accuracy due to insufficient context mixing, while performance improves as the TSP layer moves deeper. The proposed selection algorithm identifies layer 15 as an optimal point where the LongBench score begins to saturate, indicating a favorable trade-off between speed and accuracy.
TSP Length. We investigate the impact of TSP length on model performance under varying KV budgets. TSP length and KV budget control acceleration in the prefill and generation stages, respectively, and are independently tunable. As shown in Figure 8(b), accuracy drops noticeably when the TSP length is below 2048, while it saturates beyond that point. Since shorter lengths reduce latency but limit context propagation, a TSP length of 2048 provides an effective balance between latency and accuracy. This configuration is adopted in our main experiments to optimize prefill efficiency without compromising general performance.
5 Conclusion
In this paper, we present FastKV, a novel KV cache compression framework that enhances the efficiency of long-context LLM inference while preserving model accuracy. FastKV introduces Token-Selective Propagation (TSP), a mechanism that allows the early layers of the model to attend to the full input sequence, while the later layers operate only on a reduced set of semantically important tokens. This design significantly reduces the computational burden of the prefill stage without discarding critical contextual information. Additionally, FastKV compresses the KV cache for the selected tokens, enabling additional efficiency in memory usage without introducing latency bottlenecks. Experimental results show that FastKV achieves higher throughput and lower TTFT than baseline methods while maintaining high accuracy on long-context benchmarks.
Appendix A Additional Methodological Details
A.1 FastKV Algorithm
We present the pseudocode for the FastKV algorithm, which performs two complementary compression strategies during the prefill stage: (1) reducing hidden states via Token-Selective Propagation (TSP), and (2) compressing the attention key-value (KV) cache at each layer. These two steps are carried out independently: TSP selects a subset of hidden states to propagate to later layers, while KV cache compression prunes KV cache entries at each layer using attention-based importance scores. Together, they improve prefill efficiency and reduce memory usage without modifying the model architecture or introducing any runtime overhead during decoding.
The key terms used in the pseudocode are:
  • LTSP: The index of TSP layer
  • BTSP: TSP length
  • BKV: KV budget
  • KVCompress: Selects the KV entries of top-BKV critical tokens as determined by Equation 2.
  • HiddenCompress: Selects hidden states of top-BTSP critical tokens, as determined by Equation 3, to propagate to next layer.